### v1

# LEGO_HQEC: A Software Tool for Analyzing Holographic Quantum Codes

Abstract
Quantum error correction (QEC) is a crucial prerequisite for future large-scale quantum computation. Finding and analyzing new QEC codes, along with efficient decoding and fault-tolerance protocols, is central to this effort. Holographic codes are a recent class of QEC subsystem codes derived from holographic bulk/boundary dualities. In addition to exploring the physics of such dualities, these codes possess useful QEC properties such as tunable encoding rates, distance scaling competitive with topological codes, and excellent recovery thresholds. To allow for a comprehensive analysis of holographic code constructions, we introduce LEGO_HQEC, a software package utilizing the quantum LEGO formalism. This package constructs holographic codes on regular hyperbolic tilings and spawns their stabilizer generators and logical operators for a specified number of seed codes and layers. Three decoders are included: an erasure decoder based on Gaussian elimination; an integer-optimization decoder; and a tensor-network decoder. With these tools, LEGO_HQEC thus enables future systematic studies regarding the utility of holographic codes for practical quantum computing.
I Introduction
The field of quantum computing is currently focused on understanding how best to construct a large-scale fault-tolerant quantum computer [1, 2, 3, 4, 5, 6]. The physical levels of noise that dominate in current quantum computing hardware are simply too high to reliably run all but the simplest quantum algorithms. As a solution quantum error correction (QEC) techniques [7, 8, 9, 10] were proposed in order to orchestrate strategies that protect information at the logical level and facilitate the execution of a quantum algorithm on a device.
Although QEC in theory posits a method by which practically useful quantum computing could be realized, in reality the constraints imposed by current state-of-the-art code constructions and so-called fault-tolerance protocols are extremely resource-intensive [4]. As such, almost all areas of the QEC design pipeline are under active research. Some of these research areas range from the theoretical to the practical, and include: code construction [11, 12, 13]; fault tolerant error correction protocols [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]; fault-tolerant universal logical gate set implementations, such as magic-state distillation [26], gauge-fixing [27], and code switching [28, 29, 30], among others [31, 32]; decoder design [33, 34, 35, 36]; architectural proposals [1, 37, 38]; among many other engineering-level challenges [39, 12].
Holographic quantum codes are a recently introduced class of quantum error correction codes [40, 41, 13], derived as models of bulk/boundary dualities in theoretical physics, in particular the AdS/CFT correspondences [42, 43]. The main idea for these codes lies in the fact that bulk degrees of freedom, which are associated with logical qubits in $d$ dimensions, are mapped to a $(d-1)$-dimensional boundary, with the hyperbolic structure of the bulk ensuring that the physical Hilbert space is larger than its logical counterpart. Constructed on hyperbolic manifold tessellations, they can be expressed as concatenated subsystem codes and allow for extremely high (and tunable) encoding rates as well as excellent threshold performance in comparison to other leading quantum code constructions, such as quantum low-density parity-check (qLDPC) [12] and topological (TQEC) quantum codes [39, 37]. However, for all practical purposes, actually generating the conformal quasicrystal lattices [44, 45, 46] responsible for defining physical-qubit representations of stabilizers and logical operators is a challenging task, exacerbated by the need to apply suitable decoding algorithms derived from tensor-network decoding techniques [47, 48, 49, 50]. In this vein, having a software tool at hand to compute the effect of quantum noise channels is indispensable for the study of new holographic (as well as generalized concatenated) quantum codes.
In this work, we introduce a software tool named LEGO_HQEC, which has been designed to ease the code- and operator-generation process, thereby expediting and lowering the barrier to entry for this new subfield of quantum error correction. In particular, we implement a stabilizer and logical operator generator that returns a list of said objects for both maximum-rate and asymptotically zero-rate variants of many existing holographic codes, as well as providing valuable customization options for constructing homogeneous or heterogeneous constructions (i.e., holographic codes constructed from either one or several smaller seed codes) on hyperbolic manifolds parameterized by Schläfli symbols $\{p,q=4\}$ [51]. In addition to this tool, we provide three different decoders which have been used in recent studies regarding the threshold properties of holographic codes for practical quantum computing: an erasure decoder built using a Gaussian elimination (GE) method [52, 53]; a Pauli-noise integer-optimization decoder, as from [54, 52]; and a tensor-network decoder, which was utilized in [55, 48, 49] additionally for Pauli noise. With the aim of developing newer and better open-source software tools for studying quantum error correction, this package has been made publicly available at https://github.com/QML-Group/HQEC.
This paper is organized as follows: Section II gives an overview of the quantum LEGO formalism, and how it relates to stabilizer and holographic quantum error correction codes; Section III details the structure of LEGO_HQEC, starting with boundary stabilizer and logical operator generation, and ending with a concise overview of the distinct decoders offered; in Section IV, we show two examples using: the $\llbracket 5,1,3\rrbracket$ (or Harlow-Preskill-Pastawski-Yoshida (HaPPY) [40]), as well as a holographic Steane code [53]. The calculated runtime for generating stabilizers and logical operators for several layers of the asymptotically zero-rate HaPPY code is presented, as well as decoding for erasure and Pauli noise in the asymptotically zero-rate Steane code, in addition to the maximum-rate HaPPY code. We conclude by proposing several extensions to LEGO_HQEC in Section V.
II Background
II.1 Quantum Error Correction & Tensor Networks
A quantum error correction (QEC) code is a technique used to protect information at the level of physical qubits (or qudits) from deleterious environmental noise effects. Such protection is accomplished by redundantly mapping physical qubit states into subspaces of a larger, multipartite Hilbert space; this description at the level of an encoding isometry effectively defines a QEC code [9, 7].
There are many different methods by which one can construct a quantum error correction code; however, for brevity and for the scope in this article, we will focus on stabilizer quantum codes, a class of codes for which the codespace $\mathsf{C}$ is defined by the simultaneous $+1$ eigenspace of Pauli stabilizer generators; that is, operators in the codespace, which take on the following form:
$\mathsf{C}=\{\ket{\psi}\in\mathcal{H}:S\ket{\psi}=\ket{\psi},\forall S\in\mathsf{S}\}~{}.$ (1)
The set of generators $\mathsf{S}=\{S\}$ which stabilize the codespace form an Abelian subgroup which is known as the stabilizer group $\mathsf{S}$.
In order to define the logical operators of a quantum code, we first must define several terms. The normalizer $\mathcal{N}$ of the stabilizer group is defined as
$\mathcal{N}(\mathsf{S})=\{\mathsf{P}\in\mathsf{P}_{n}:\mathsf{P}\mathsf{S}\mathsf{P}^{\dagger}=\mathsf{S}\},$ (2)
where $\mathsf{P}$ is an element of the $n$-fold Pauli group $\mathsf{P}_{n}$. Then, we define the logical operators of a quantum code as pertaining to the group $\mathcal{N}(\mathsf{S})/\mathsf{S}$.
Subsystem codes are typically stabilizer codes in which the encoded subspace is split up and subdivided. Normally, the Hilbert space of a stabilizer code is divided as $\mathcal{H}=\mathcal{H}_{C}\oplus\mathcal{H}_{\bar{C}}$, where $\mathcal{H}_{C}$ is used to store the logical information, and $\mathcal{H}_{\bar{C}}$ is the complement. We can additionally subdivide $\mathcal{H}_{C}$ further as $\mathcal{H}_{C}=\mathcal{H}_{L}\otimes\mathcal{H}_{J}$, where $\mathcal{H}_{L}$ is the logical subspace and $\mathcal{H}_{J}$ is known as the junk (or gauge) subspace [7]. The gauge subspace is used to help diagnose and correct errors inside of $\mathcal{H}_{L}$; such a feat is accomplished by defining the gauge group such that the stabilizer subgroup is the center $\mathsf{Z}(\mathsf{G})$ of the gauge group
$\mathsf{Z}(\mathsf{G}):=\{i\mathcal{I},\mathsf{S}\}=\mathsf{C}(\mathsf{G})\cap\mathsf{G}~{},$ (3)
where $\mathsf{G}$ is the junk (or gauge) group associated with $\mathcal{H}_{J}$, and $\mathsf{C}(\mathsf{G})$ is the centralizer of $\mathsf{G}$, which is defined as $\mathsf{C}(\mathsf{G}):=\{\mathcal{O}\in\mathsf{G}|[\mathcal{O},S]=0,\forall S\in\mathsf{S}\}$.
As a specific example of a stabilizer code, we introduce here the Steane code [56, 57], which is a Calderbank-Shor-Steane code. This $\llbracket 7,1,3\rrbracket$ code has X and Z stabilizers symmetrically split in the parity check matrix, and the stabilizer generators take the form
$X_{1}X_{4}X_{6}X_{7}~{},$ (4)
$X_{2}X_{4}X_{5}X_{7}~{},$ (5)
$X_{3}X_{5}X_{6}X_{7}~{},$ (6)
$Z_{1}Z_{4}Z_{6}Z_{7}~{},$ (7)
$Z_{2}Z_{4}Z_{5}Z_{7}~{},$ (8)
$Z_{3}Z_{5}Z_{6}Z_{7}~{},$ (9)
with logical operators defined as
$\bar{X}$ $=X_{1}X_{2}X_{3}X_{4}X_{5}X_{6}X_{7}\ ,$ (10)
$\bar{Z}$ $=Z_{1}Z_{2}Z_{3}Z_{4}Z_{5}Z_{6}Z_{7}\ .$ (11)
Another small example of a quantum error correction code is the perfect $\llbracket 5,1,3\rrbracket$ code, which is known as the smallest $d=3$ code [58]; its stabilizer generators are
$X_{1}Z_{2}Z_{3}X_{4}~{},$ (12)
$X_{2}Z_{3}Z_{4}X_{5}~{},$ (13)
$X_{1}X_{3}Z_{4}Z_{5}~{},$ (14)
$Z_{1}X_{2}X_{4}Z_{5}~{},$ (15)
with the logical operators defined as
$\bar{X}$ $=X_{1}X_{2}X_{3}X_{4}X_{5}~{},$ (16)
$\bar{Z}$ $=Z_{1}Z_{2}Z_{3}Z_{4}Z_{5}~{}.$ (17)
These two examples in particular are important, as they constitute some of the elemental seed codes for constructing a holographic quantum code via the quantum LEGO (qLEGO) formalism [49, 59, 60]. More succinctly, the qLEGO formalism provides a method by which tensor networks can be employed in order to construct and simulate large QEC codes from seed codes. One starts by defining a quantum state as
|$\ket{\psi}=T_{i_{1}\cdots i_{m}}\ket{i_{1}\cdots i_{m}}~{},$ (18)
where the tensor $T_{i_{1}\cdots i_{m}}$ also can be used to describe a mapping from input subsystems to output subsystems, as shown in (a), (b) of Figure 1. Here the repeated indices are summed over. If qubits $1$ through $k$ are maximally entangled, we can write an isometry for $T_{i_{1}\cdots i_{m}}$ as
$V_{\ket{\bar{\psi}}}:=T_{i_{1}\cdots i_{m}}\ket{i_{k+1}\cdots i_{m}}\bra{i_{1}\cdots i_{k}}~{},$ (19)
in which we can easily take the incoming $i_{1}\cdots i_{k}$ input legs as the encoding qudits for an $\llbracket n,k\rrbracket$ quantum error correction code (where $n,k$ represent the numbers of physical and logically-encoded qudits, respectively), following [59]. In general, it is not necessary that $T_{i_{1}\cdots i_{m}}$ represent an isometry [59]; however, a convenient choice is to require that $T_{i_{1}\cdots i_{m}}$ corresponds to a multipartite maximally-entangled (MME) state [61], thereby necessitating that a many-body quantum state adhere to the condition
$\rho_{s}=\text{Tr}_{s^{c}}\big{[}\ket{\psi}\bra{\psi}\big{]}\propto\mathbb{I}~{},$ (20)
where $s\in\{1\cdots n\}$, $|s|\leq\left\lfloor{\frac{n}{2}}\right\rfloor$, and $s^{c}$ denotes the complementary set of subsystems to $s$. States adhering to this property are known as k-uniform, and have applications in many areas of quantum information theory [61]. Note that the qLEGO formalism also applies to more general tensors that do not form isometries, the contraction of isometric tensors (such as those above) permit the straightforward formation of a quantum code’s encoding circuit [59].
Continuing, we find a graphical description in Figure 1(c)-(d) of the concept known as tensor pushing. Simply put, since $T_{i_{1}\cdots i_{m}}$ can form an isometry from $i_{1}\cdots i_{k}$ to $i_{k+1}\cdots i_{m}$, it is possible to input an operator $\mathcal{O}$ with support on the input legs (as shown) and push the operator past the isometry, finding a new representation of the operator on the output legs.
Figure 1: Tensor-network diagrams depicting (a)-(b) the tensor $T_{i_{1}\cdots i_{m}}$ as it pertains to a quantum state, as well as an isometric map via the Choi-Jamiolkowski isomorphism [8], and (c)-(d) the action of tensor pushing an operator $\mathcal{O}$ through the input subsystems of the isometry in order to find an output-leg representation, which we name $\mathcal{O}^{\prime}$.
Figure 2(a) shows the support of a unitary logical operator $\mathcal{\bar{O}}$ on the output subsystems as $\mathcal{\bar{O}}=\mathcal{O}_{1}\otimes\cdots\otimes\mathcal{O}_{6}$ (in our specific case, for simplicity). In (b), we see that contracting two such tensors along their respective indices gives rise to a new conjoined tensor (in this case with two logical qubits). Finding the new logical operators for this new isometry is quite simple, and follows the tensor-matching formalism described in detail in [59, 60]. As shown in (c), the principle requirement is that support on contracted output indices must be identical up to complex conjugation; in the specific case of Figure 2(c), the requirements must be that $\mathcal{O}_{3}=\mathcal{O}^{\prime*}_{6}$ and $\mathcal{O}_{4}=\mathcal{O}^{\prime*}_{5}$.
Figure 2: (a) One may define an operator $\mathcal{\bar{O}}$ on the input legs to a tensor by its representation (in the present case $\mathcal{O}_{1}\otimes\cdots\otimes\mathcal{O}_{6}$) on the output legs. (b) Contracting indices together yields a new, conjoined tensor network. (c) Deducing the new logical operator is possible by keeping the original logical operators that satisfy the matching condition $\mathcal{O}_{3}=\mathcal{O}^{\prime*}_{6}$ and $\mathcal{O}_{4}=\mathcal{O}^{\prime*}_{5}$ [59, 60].
II.2 Holographic Quantum Codes
Holographic quantum codes are a relatively new class of concatenated subsystem codes [62, 52, 63]. These codes have been found to exhibit properties consistent with the famous AdS/CFT correspondence [41, 43, 42, 64], in addition to possessing interesting and potentially useful error-correction properties for practical quantum computing, such as: high, tunable rates [52, 48]; versatility in constructions involving several different seed codes [62]; distance scaling better than $\sqrt{n}$ [54]; and constructions with large fault-tolerant logical gate sets [62], among other useful properties [55, 65].
Figure 3: Examples of several holographic tensor-network codes that LEGO_HQEC can construct and test. (a) The maximum-rate Harlow-Preskill-Pastawski-Yoshida (HaPPY) code on a hyperbolic pentagon tiling, mapping from logical (red) to physical (black) legs (Section IV.1.1). (b) A partially gauge-fixed version of the HaPPY code is created by projecting some logical legs onto an eigenstate |$\ket{G}$ of a gauge operator $G$. (c) The asymptotically zero-rate holographic Steane code, which uses the same 8-leg tensor once as a logical-to-physical map (central heptagon tensor) and otherwise with only planar legs (octagon tensors); a test of this code is performed in Section IV.1.2.
A holographic quantum code is built typically from either one or several seed codes, as shown in Figure 3. As hyperbolic tessellations allow for many different orientations and polygonal embedments in the bulk according to the famous Schläfli symbol formula $\frac{1}{p}+\frac{1}{q}<\frac{1}{2}$, infinitely many variations of holographic quantum codes are possible to construct [62, 51]. However, only homogeneous [53, 40, 49, 52] (that is, those holographic codes constructed from uniform hyperbolic tessellations with only one polygon featured, as shown in Figure 3(a)) and simple heterogeneous (with only alternating layers of two seed codes, as in Figure 3(b)) holographic codes [62] have so far been constructed. Growing a holographic code layerwise is done in two distinct methods, as shown in Figure 4: either using (a) edge inflation, in which every edge is surrounded by vertices in the forthcoming layers of the tensor network, or (b) vertex inflation, in which each vertex is surrounded by edges in each subsequent layer. In the tensor-network scheme, the concatenation is apparent since subsequent layers must contract indices of the previous layer in order to adhere to the stringent constraints of a hyperbolic tessellation. A diagrammatic representation of these two methods can be seen in Figure 4. More details on this topic and in general on holographic codes can be found in [13].
Figure 4: Edge inflation (a) and vertex inflation (b) of a tensor network on the $\{5,4\}$ face-centered hyperbolic tessellation [44, 45, 13]. In both cases, we show the first three inflation steps around a central pentagon.
III Software Design
LEGO_HQEC was developed in Python, with ease of use front and center in the design considerations. The overarching design premise of LEGO_HQEC involves automation and extendability. As such, LEGO_HQEC is modularized into two parts: (a) the stabilizer and logical operator generator, and (b) the decoders. Both will be addressed in what follows.
III.1 Stabilizer & Logical Operator Generation
Stabilizer and logical operator generation is carried out following several steps, which are depicted in Figure 5. Firstly, the object TensorNetwork is defined, with dependencies Tensor and TensorLeg as important classes; these are the components used to construct the holographic tensor network. Next, the Push function is defined; in this portion, each tensor has its own OperatorPush method on its local legs. Push then activates tensors one-by-one from the center to the boundary, and sequentially generates boundary representations of all locally-defined stabilizers and logical operators. Additionally, legs that have undergone an operator push will be blocked, in order to prevent operators from being accidentally pushed back in towards the center; in this way a locally-defined operator can be pushed to the boundary of the network. Finally, the ReadOut layer retrieves boundary stabilizers and logical operators using a backtracking method, and compiles them into the final boundary result; this backtracking method starts from the center of the tensor network, then uses a depth-first traversal to reach outer tensors, while recording operators on the legs of the outermost tensors at the boundary. With this method, it is ensured that boundary operators align with geometric intuition from holographic codes. A a last step, a correctness check is performed by again traversing all internal legs and checking to see that all bulk-defined operators have been pushed to the boundary. More details on this portion of the software are available at [65].
Figure 5: Pictographic representation of the stabilizer / logical operator generator in LEGO_HQEC. TensorNetwork is a list containing all of the Tensor objects, whose attributes include the TensorLeg object. This list of tensors is then filtered into the Push function, which elementwise processes and pushes each operator from each tensor to the boundary, obtaining a boundary representation. Finally, a correctness checker (CorrectnessChecker) backtracks along the tensor network, checking to see that each contracted index from each tensor contains only the identity operator. The final boundary representation is read out with ReadOut.
III.2 Decoders
Three decoders are included in LEGO_HQEC: an erasure decoder (Section III.2.1) for the quantum erasure channel [8, 9], as well as two decoders for Pauli noise based on integer optimization and a tensor-network decoder (Section III.2.2) [48, 49].
III.2.1 Quantum Erasure Channel
Erasure (or qubit loss) is a quantum noise channel $\mathcal{E}_{e}(\cdot)$ with two inputs (|$\ket{0},\ket{1}$) and three outputs (|$\ket{0},\ket{1},\ket{e}$) [8, 66, 9], and can be written as
$\mathcal{E}_{e}(\rho)=(1-\epsilon)\rho+\epsilon\text{Tr}\big{[}\rho\big{]}\ket{e}\bra{e}~{},$ (21)
where with some probability $(1-\epsilon)$ a qubit is left unchanged and with probability $\epsilon$ the qubit is “erased”. |$\ket{e}$ is a state orthogonal to the input state $\rho$.
LEGO_HQEC includes an erasure decoder based on Gaussian elimination (GE), as was utilized in [52, 53]. The general idea of the GE decoder involves stacking all of the stabilizers and logical operators into a large matrix, and then using the support of the erasure vector, for which entries are randomly generated per qubit with probability $p$, to filter out columns that have undergone an erasure. The algorithm then proceeds by executing Gaussian elimination to find the reduced-row-echelon form of the matrix; if it is found that both logical operators still have physical support on the remaining qubits, then the trial is counted as a success. The runtime complexity of the GE algorithm is known to scale as $\mathcal{O}(n(n-k)^{2})$. More details on the GE decoder can be found at [53, 52].
III.2.2 Pauli Noise Channels
The prototypical Pauli noise channel is represented by the depolarizing noise channel $\mathcal{E}_{d}(\cdot)$ [9, 8], and takes the form
$\mathcal{E}_{d}(\rho)=(1-\delta)\rho+\frac{\delta}{3}\big{(}X\rho X+Y\rho Y+Z\rho Z\big{)}~{},$ (22)
where $\delta$ is the probability of an error occurring, and $\{X,Y,Z\}$ represent the three Pauli operators. For Pauli noise, LEGO_HQEC employs two different types of decoders: one based on integer optimization [52, 54], and another based on tensor-network decoding [55, 49].
Integer optimization decoding works by a very similar technique as for the GE decoder discussed in Section III.2.1, but with some small changes. Firstly, as error locations are not known, one must first calculate the Moore-Penrose pseudoinverse [67, 68], which corresponds to the destabilizers (or pure errors) of the code [54, 69]. One then obtains a guessed error by multiplying the pseudoinverse matrix directly by the vector of the obtained syndrome. Finally, an integer optimization is utilized in order to deduce the lowest-weight pure-error correction [65, 54]. It is known that integer-optimization decoding exhibits with a runtime complexity of $\mathcal{O}(2^{n})$, which is dependent on the size of the quantum code. More details on integer-optimization decoding can be found in [65, 52, 54].
Tensor-network decoding, however, works by exploiting tensor-network geometry and clever contraction sequences [49, 48]. By setting the bond dimension of each tensor in the network to be four, Pauli errors $\{I,X,Y,Z\}$ can be easily accounted for. In this way, an error vector is initialized at random with some error probability $d$. These error probabilities are then utilized in the tensor network, and, upon contraction, yield a close approximation in the thermodynamic limit to maximum-likelihood decoding [50]. It is known that tensor-network decoding exhibits a runtime complexity of $\mathcal{O}(n^{2.37})$ in the best-case scenario [70, 71], and that multiple logical qubits can be decoded in parallel as $\mathcal{O}(nk)$, as long as the error rate stays below the critical error rate of the central logical qubit (i.e., the threshold of the central logical qubit) [48]. More details on tensor-network decoding can be found in [47, 49, 48].
IV Examples Using LEGO_HQEC
This section contains a step-by-step set of examples for utilizing LEGO_HQEC. We begin by initializing and testing a maximum-rate Harlow-Preskill-Pastawski-Yoshida (HaPPY) [40] code under erasure noise, and then following up with an asymptotically zero-rate Steane code [53] under Pauli depolarizing noise. The results of both decoder simulations are shown in Figure 6.
Figure 6: (a) Erasure decoder results for the maximum-rate HaPPY code (which obtains no threshold, as per the results in [40]), which utilized the GE decoder (Section III.2.1). (b) Pauli depolarizing noise results from the tensor-network decoder for an asymptotically zero-rate Steane code; the threshold of this code is found to be approximately $18.25\%$, in agreement with results from [54]. $750$ and $500$ Monte Carlo trials were utilized for the erasure and depolarizing noise simulations, respectively. Results are gathered per layer of concatenation $L$.
IV.1 Usage
IV.1.1 Maximum-Rate HaPPY Code under Erasure Noise
1from OperatorPush.PushingToolbox import batch_push, batch_push_multiprocessing
2from QuDec.InputProcessor import extract_stabilizers_from_result_dict, extract_logicals_from_result_dict
3from QuDec.OutputProcessor import save_results_to_csv
4import numpy as np
5from QuDec.ErasureDecoder import calculate_recovery_rates_for_p_range
6from OperatorPush.Presets.Zero_Rate_HaPPY_new_for_kj import setup_zero_rate_happy, setup_max_rate_happy
7if __name__ == ’__main__’:
8 # HaPPY Code, max rate
9 for R in [0, 1, 2]:
10 tensor_list = setup_max_rate_happy(R=R)
11 results_dict = batch_push_multiprocessing(tensor_list)
12 stabilizers = extract_stabilizers_from_result_dict(results_dict)
13 logical_zs, logical_xs = extract_logicals_from_result_dict(results_dict)
14 print("logical_xs: ", logical_xs, logical_zs)
15 logical_operators = [logical_zs[0]] + [logical_xs[0]]
16 results = calculate_recovery_rates_for_p_range(n=750, p_start=0.01, p_end=1.0, p_step=0.01, stabilizers=stabilizers,logical_operators=logical_operators,n_process=1, cpu_affinity_list=None)
17 save_results_to_csv(results, file_path=F’R{R}_rec_happy.csv’)
Listing 1: Code snippet for initializing the stabilizers and logical operators of the maximum-rate HaPPY code, as well as for testing the first three layers of concatenation under the erasure noise channel. Decoding was carried out using the GE decoder.
LABEL:code1 displays a code snippet for obtaining the threshold of the maximum-rate HaPPY code under erasure noise, utilizing the GE decoder discussed in Section III.2. In the example above, the commensurate libraries and dependencies are first imported (lines 1-6), and then the stabilizers and logical operators for each layer are generated (lines 10-13). Next, those objects are then fed into the erasure decoder per layer $L$, and the Monte Carlo trials commence layerwise (lines 9 and 16). Finally, the results are written into a csv file and saved (lines 19).
IV.1.2 Asymptotically Zero-Rate Steane Code under Depolarizing Noise
1from QuDec.TN_decoder import tn_quantum_error_correction_decoder_multiprocess
2import numpy as np
3from OperatorPush.PushingToolbox import batch_push, batch_push_multiprocessing
4from QuDec.InputProcessor import extract_stabilizers_from_result_dict, extract_logicals_from_result_dict
5from QuDec.OutputProcessor import save_results_to_csv
6from OperatorPush.Presets.Heptagon_Steane_Code import setup_heptagon_zero_rate_steane
7if __name__ == ’__main__’:
8 # Steane Code, zero rate
9 for R in [0, 1, 2]:
10 p_depo_step = 0.01
11 p_depo_start = 0.1
12 p_depo_end = 0.5
13
14 tensor_list = setup_heptagon_zero_rate_steane(R=R)
15 results_dict = batch_push_multiprocessing(tensor_list)
16 stabilizers = extract_stabilizers_from_result_dict(results_dict)
17 logical_zs, logical_xs = extract_logicals_from_result_dict(results_dict)
18 success_rates = []
19 rx = 0.33
20 rz = 0.33
21 ry = 1 - rx - rz
22 for p in np.arange(p_depo_start, p_depo_end, p_depo_step):
23 results = tn_quantum_error_correction_decoder_multiprocess(tensor_list=tensor_list, p=p, N=500, rx=rx, ry=ry, rz=rz, stabilizers=stabilizers, logical_z=logical_zs[0], logical_x=logical_xs[0], n_process=1)
24 success_rates.append((p, results))
[…]
Listing 2: Code snippet for initializing the stabilizers and logical operators of the asymptotically zero-rate Steane code, as well as for testing the first three layers of concatenation under the Pauli depolarizing noise channel using the tensor-network decoder.
LABEL:code2 shows the code needed for initializing the tensor-network decoder for a depolarizing noise channel. In contrast to the previous case, the tensor network decoder requires a specification of the relative proportions of X, Y, and Z errors (lines 19-21). However, the actual generation of the stabilizers and logical operators proceeds identically as for the HaPPY code (lines 14-17). In the future, functionality will be added in order to allow the user to specify which tensors will have usable logical qubits, as well as gauge-fixing patterns that have featured in other holographic quantum codes recently [52, 63].
IV.2 Runtime Performance
Here we report an example of the runtime performance expected of an asymptotically zero-rate HaPPY code for: the stabilizer and logical operator generation OperatorPush; the GE erasure decoder; the integer-optimization decoder; and for the tensor-network decoder. All results are given as a function of layer $L$ and times are reported in seconds (s). For the decoders, the plotted results are per Monte Carlo trial, while the OperatorPush results denote the per layer start-to-finish time for pushing all locally-defined stabilizers and logical operators to their respective boundary representations, and extracting them into a csv file.
Figure 7: Time-scaling results per event for the OperatorPush function, as well as the three different decoders utilized in this software package. Due to limitations in computational power, we show the results regarding the exponentially-scaling integer-optimization algorithm for only the first two layers of the asymptotically zero-rate HaPPY code.
As we were limited by computational power, we were able to extract results for only the first two layers using the integer-optimization decoder. For all other decoders and the OperatorPush method, however, we were able to obtain time-scaling estimates for up to $L=4$, i.e. on the order of roughly several thousand boundary (physical) qubits, depending on the Schläfli symbols $\{p,q\}$ associated with the commensurate hyperbolic tiling.
V Conclusion
In this work, we have presented and demonstrated LEGO_HQEC, a new software tool for constructing and testing holographic quantum codes under typical noise channels of interest. Similar to other notable software packages for quantum error correction, fault tolerance, among others [72, 73, 74, 75, 76, 77], our focus with LEGO_HQEC has been to facilitate open-source development and enable future scientific discovery in the realm of holographic quantum error correction. In the process, we have: developed and implemented an easy-to-use open-source software tool; demonstrated several examples of its capabilities; and have provided a basic framework upon which more detailed code studies and further discoveries in holographic quantum error correction can be leveraged.
LEGO_HQEC has already featured in several holographic code studies, including those exploring the novel Evenbly codes [52, 63], analyses of asymptotically zero-rate holographic codes as compared to the hashing bound [55, 65], as well as for constructions of heterogeneous holographic codes [62], which are constructed from several seed codes and can circumvent the Eastin-Knill theorem regarding codes with universal logical gate sets [78, 31, 32].
In [49, 60, 76], tensor-network decoders have been implemented using the Julia language (which is available at https://github.com/qecsim/TensorNetworkCodes.jl). Although one can technically build holographic tensor-network codes using this package, the principal difference lies in the fact that our work is streamlined and specialized for generating regular hyperbolic lattices with a quasiregular boundary in an efficient manner, as well as for testing codes under a wide range of possible noise channels; these include the erasure, depolarizing, and biased-noise channels (such as those explored in [55]).
There are many ideas for future modification of LEGO_HQEC. Firstly, an extension for considering any hyperbolic tiling parameterized by $\{p,q\}$ would be enormously helpful, as it has been shown that certain classes of holographic codes have vastly different threshold behavior as the tiling is altered [52]. Furthermore, recent work [62, 65] shows that incorporating multiple seed codes can allow for potentially universal fault-tolerant logical operations [62]; tools to support such fault-tolerance studies could be very useful. One can also consider implementing functionality for greater control over the encoded logical qubits in the bulk, as recent work shows that recovery rates of finite-rate holographic codes greatly depend not only on the number $k$ of encoded logical qubits, but critically on the positions of said logical qubits in the bulk [79]. Even more broadly, it is known that the qLEGO formalism applies to non-Abelian stabilizer codes, such as the XP stabilizer variety [80, 81], or even non-stabilizer codes, such as those presented in [82, 7, 83, 84, 85]; one may consider incorporating functionality to accommodate such holographic code constructions in the near future. Moreover, the class of codes known as Evenbly codes, a derivate of the hyperinvariant tensor networks first introduced in [86, 87], have been extensively studied using LEGO_HQEC [52, 63]; however, one major impediment was the inclusion of Hadamard gates on the contracted edges of the tensor network, thereby hindering decoding of Pauli channels with the tensor-network decoder. Allowing for such constraints, among others, remains within the future goals of this work. Finally, there are many additions purely from the software and usability standpoint, such as: adding functionality for tree-style concatenated codes [23, 10, 7]; visualization tools; or GPU acceleration tools; among other prospects.
V.1 Acknowledgements
We thank David Elkouss and Rob Harris for useful discussions. MS and SF acknowledge financial support from the Intel corporation. AJ is supported by the Einstein Research Unit “Perspectives of a quantum digital transformation”. CC acknowledges support from the Commonwealth Cyber Initiative at Virginia Tech.
V.2 Author Contributions
JF and MS implemented the Gaussian elimination decoder. JF developed the integer-optimization decoder, tensor-network decoder, and the OperatorPush generator during his master thesis; the master thesis itself was supervised by MS and SF. AS provided assistance with software design. MS wrote the manuscript, with the hyperbolic tiling figures provided by AJ. CC, AJ, AS, and SF provided project guidance and supervision.