### v3

# and Decoding of Holographic Quantum Codes

Abstract
Quantum error correction (QEC) is a crucial prerequisite for future large-scale quantum computation. Finding and analyzing new QEC codes, along with efficient decoding and fault-tolerance protocols, is central to this effort. Holographic codes are a recent class of generalized concatenated codes derived from holographic bulk/boundary dualities. In addition to exploring the physics of such dualities, these codes possess useful QEC properties such as tunable encoding rates, distance scaling competitive with other well-studied code classes, and excellent recovery thresholds. To allow for a comprehensive analysis of holographic code constructions, we introduce LEGO_HQEC, a software package utilizing the quantum LEGO formalism. This package allows for the construction and analysis of holographic codes on regular hyperbolic tilings, computing their stabilizer generators and logical operators for a specified number of seed codes and layers. Three decoders are included: an erasure decoder based on Gaussian elimination; an integer-optimization decoder; and a tensor-network decoder. With these tools, LEGO_HQEC enables systematic studies of both previously known holographic codes and novel variants. As a demonstration, we provide new numerical results on the holographic black-hole pentagon code, establishing its threshold behavior under the erasure channel as a benchmark example.
I Introduction
The field of quantum computing is currently focused on understanding how best to construct a large-scale fault-tolerant quantum computer [campbell2017roadsterhal, gottesman2013fault, preskill2018quantum]. The physical levels of noise that dominate in current quantum computing hardware are simply too high to reliably run all but the simplest quantum algorithms. As a solution, quantum error correction (QEC) techniques [lidar2013quantum, wilde2013quantum, nielsen2010quantum, gottesman1997stabilizer] were proposed in order to orchestrate strategies that protect information at the logical level and facilitate the reliable execution of quantum algorithms on quantum devices.
Although QEC in theory posits a method by which practically useful quantum computing could be realized, in reality the constraints imposed by current state-of-the-art code constructions and so-called fault-tolerance protocols are extremely resource-intensive [gottesmanFT]. As such, almost all areas of the QEC design pipeline are under active research. These research areas range from the theoretical to the practical, and include: code constructions [albert2022bosonic, breuckmann2021quantum, jahn_topical]; fault tolerant error correction protocols [between_shorsteane, dhruvi_groovy, chamberland2018flag, delfosse2021beyond, bombin2015single, divincenzo1996fault, divincenzo2007effective, chao2018quantum, aliferis2005quantum, knill2005quantum, shor1996fault, delfosse2020short]; fault-tolerant universal logical gate set implementations, such as magic-state distillation [bravyi2005universal], gauge-fixing [bombin2015gauge], and code switching [codeswitching1, codeswitching2, codeswitching3], among others [jochym2014using, chamberland2016thresholds]; decoder design [iolius2023decoding, dinur2023good, roffe2020decoding, babar2015fifteen]; architectural proposals [campbell2017roadsterhal, terhal2015quantum, yoshida2024concatenate]; among many other engineering-level challenges [fowler2012surface, breuckmann2021quantum].
Holographic quantum codes are a recently introduced class of quantum error correction codes [preskill_happy_code, harlow2015bulk, jahn_topical], derived as models of bulk/boundary dualities in theoretical physics, in particular the AdS/CFT correspondences [maldacena1999large, witten1998anti]. The main idea for these codes lies in the fact that bulk degrees of freedom, which are associated with logical qubits in $d$ dimensions, are mapped to a (d1)(d-1)-dimensional boundary, with the hyperbolic structure of the bulk ensuring that the physical Hilbert space is larger than its logical counterpart. Constructed on hyperbolic manifold tessellations, they can be expressed as generalized concatenated codes and allow for extremely high (and tunable) encoding rates as well as excellent threshold performance in comparison to other leading quantum code constructions, such as quantum low-density parity-check (qLDPC) [breuckmann2021quantum] and topological (TQEC) quantum codes [fowler2012surface, terhal2015quantum]. However, for all practical purposes, actually generating the conformal quasicrystal lattices [conformal_quasicrystals_holography, jahn2022tensor_qcft, central_charge_aperiodic] responsible for defining physical-qubit representations of stabilizers and logical operators is a challenging task, exacerbated by the need to apply suitable decoding algorithms derived from tensor-network decoding techniques [joe_chubb, farrelly_parallel_decoding, farrelly_tn_codes, chubb2021statistical]. In this vein, having a software tool at hand to compute the effect of quantum noise channels is indispensable for the study of new holographic (as well as generalized concatenated) quantum codes.
In previous studies [farrelly_tn_codes, Farrelly_ltnc, qecsim], tensor-network decoders have been implemented using the Julia language (which is available at https://github.com/qecsim/TensorNetworkCodes.jl). Although one can in principle also build holographic tensor-network codes using this package, our work provides a specific advantage for studying holographic codes due to its streamlined setup and its specialization to generating regular hyperbolic lattices with a quasiperiodic boundary in an efficient manner, as well as its ability to test codes under a wide range of possible noise channels; these include the erasure, depolarizing, and biased-noise channels (such as those explored in [fan2024overcoming]).
In this manuscript, we introduce a novel software tool named LEGO_HQEC in order to expedite and lower the barrier to entry for this new subfield of quantum error correction. In particular, we implement a stabilizer and logical operator generator that returns a list of these for both maximum-rate and asymptotically zero-rate variants of many existing holographic codes, as well as providing valuable customization options for constructing homogeneous or heterogeneous constructions (i.e., holographic codes constructed from either one or several smaller seed codes) on hyperbolic manifolds parameterized by Schläfli symbols $\{p,q{=}4\}$ [ratcliffe1994foundations]. In addition to this tool, we provide three different decoders which have been used in recent studies regarding the threshold properties of holographic codes for practical quantum computing: an erasure decoder built using a Gaussian elimination (GE) method [steinberg2024far, harris_css_hqec]; a Pauli-noise integer-optimization decoder, as from [harris_int_opt, steinberg2024far]; and a tensor-network decoder, which was utilized in [fan2024overcoming, farrelly_parallel_decoding, farrelly_tn_codes] additionally for Pauli noise. With the aim of developing newer and better open-source software tools for studying quantum error correction, this package has been made publicly available at https://github.com/QML-Group/HQEC. In this way, the overarching design premise of LEGO_HQEC involves automation and extendability for creating, analyzing, and decoding holographic quantum codes.
This paper is organized as follows: Section II gives an overview of the quantum LEGO formalism, and how it relates to stabilizer and holographic QEC codes; Section III details the structure of LEGO_HQEC, starting with boundary stabilizer and logical operator generation, and ending with a concise overview of the distinct decoders offered; in Section IV, we show two examples using: the $\llbracket 5,1,3\rrbracket$ pentagon (or Harlow-Preskill-Pastawski-Yoshida (HaPPY) [preskill_happy_code]) code, as well as a holographic Steane code [harris_css_hqec]. We further present novel numerical results on the black-hole HaPPY code in Section IV-A1, highlighting its threshold behavior under the erasure channel as a demonstration of the tool’s utility. The calculated runtime for generating stabilizers and logical operators for several layers of the asymptotically zero-rate HaPPY code is presented, as well as decoding for erasure and Pauli noise in the asymptotically zero-rate Steane code, in addition to the maximum-rate HaPPY code. We conclude by proposing several extensions to LEGO_HQEC in Section VI.
II Background
II-A Quantum Error Correction & Tensor Networks
Quantum error correction (QEC) codes are used to protect information at the level of physical qubits (or qudits) from deleterious environmental noise effects. Such protection is accomplished by redundantly mapping physical qubit states into subspaces of a larger, multipartite Hilbert space; this description at the level of an encoding isometry effectively defines a QEC code [nielsen2010quantum, lidar2013quantum].
There are many different methods by which one can construct a QEC code; however, for brevity and for the scope in this article, we will focus on quantum stabilizer codes, a class of codes for which the codespace $\mathsf{C}$ is defined by the simultaneous +1+1 eigenspace of Pauli stabilizer generators; that is,
[…]
wherein the set of generators $\{S\}$ which stabilize the codespace form an Abelian subgroup which is known as the stabilizer group $\mathsf{S}$, and as such $\{S\}\in\mathsf{S}$. In order to define the logical operators of a quantum code, we first must define several terms. The normalizer $\mathcal{N}$ of the stabilizer group is defined as
$\mathcal{N}(\mathsf{S})=\{\mathsf{P}\in\mathsf{P}_{n}:\mathsf{P}\mathsf{S}\mathsf{P}^{\dagger}=\mathsf{S}\},$ (2)
where $\mathsf{P}$ is an element of the $n$-fold Pauli group $\mathsf{P}_{n}$. Then, we define the logical operators of a quantum code as pertaining to the group $\mathcal{N}(\mathsf{S})/\mathsf{S}$.
As a specific example of a stabilizer code, we introduce here the Steane code [steanecode1, steanecode2], which belongs to the class of Calderbank-Shor-Steane codes. This $\llbracket 7,1,3\rrbracket$ code has $X$ and $Z$ stabilizers symmetrically split in the parity check matrix, and the stabilizer generators take the form
$X_{1}X_{4}X_{6}X_{7}~,$ (3a)
$X_{2}X_{4}X_{5}X_{7}~,$ (3b)
$X_{3}X_{5}X_{6}X_{7}~,$ (3c)
$Z_{1}Z_{4}Z_{6}Z_{7}~,$ (3d)
$Z_{2}Z_{4}Z_{5}Z_{7}~,$ (3e)
$Z_{3}Z_{5}Z_{6}Z_{7}~,$ (3f)
with logical operators defined as
$\bar{X}$ $=X_{1}X_{2}X_{3}X_{4}X_{5}X_{6}X_{7}\ ,$ (4a)
$\bar{Z}$ $=Z_{1}Z_{2}Z_{3}Z_{4}Z_{5}Z_{6}Z_{7}\ .$ (4b)
Another small example of a QEC code is the perfect $\llbracket 5,1,3\rrbracket$ code, which is known as the smallest code with code distance $d=3$ [laflamme1996perfect]; its stabilizer generators are
$X_{1}Z_{2}Z_{3}X_{4}~,$ (5a)
$X_{2}Z_{3}Z_{4}X_{5}~,$ (5b)
$X_{1}X_{3}Z_{4}Z_{5}~,$ (5c)
$Z_{1}X_{2}X_{4}Z_{5}~,$ (5d)
with the logical operators defined as
$\bar{X}$ $=X_{1}X_{2}X_{3}X_{4}X_{5}~,$ (6a)
$\bar{Z}$ $=Z_{1}Z_{2}Z_{3}Z_{4}Z_{5}~.$ (6b)
These two examples in particular are important, as they constitute some of the elemental seed codes for constructing a holographic quantum code via the quantum LEGO (qLEGO) formalism [farrelly_tn_codes, quantumlego, Farrelly_ltnc]. More succinctly, the qLEGO formalism provides a method by which tensor networks can be employed in order to construct and simulate large QEC codes from seed codes. One starts by defining a quantum state as
[…]
where the tensor $T_{i_{1}\cdots i_{m}}$ also can be used to describe a mapping from input subsystems to output subsystems, as shown in Figure 1(a)-(b). Here the repeated indices are summed over. If qubits $1$ through $k$ are maximally entangled, we can write an isometry for $T_{i_{1}\cdots i_{m}}$ as
[…]
in which we can easily take the incoming $i_{1}\cdots i_{k}$ input legs as the encoding qudits for an $\llbracket n,k\rrbracket$ QEC code (where $n$ and $k$ represent the number of physical and logically-encoded qudits, respectively), following [quantumlego]. In general, it is not necessary that $T_{i_{1}\cdots i_{m}}$ represent an isometry [quantumlego]; however, a convenient choice is to require that $T_{i_{1}\cdots i_{m}}$ corresponds to a multipartite maximally-entangled (MME) state [zyczkowski_mme], thereby necessitating that a many-body quantum state adheres to the condition
[…]
where $s\in\{1\cdots n\}$, $|s|\leq\left\lfloor{\frac{n}{2}}\right\rfloor$, and $s^{c}$ denotes the subsystems set complementary to $s$. States adhering to this property are known as k-uniform, and have applications in many areas of quantum information theory [zyczkowski_mme]. Note that the qLEGO formalism also applies to more general tensors that do not form isometries, as the contraction of isometric tensors (such as those above) permits the straightforward formation of a quantum code’s encoding circuit [quantumlego].
Continuing, we find a graphical description in Figure 1(c)-(d) of the concept known as tensor pushing. Simply put, since $T_{i_{1}\cdots i_{m}}$ can form an isometry from $i_{1}\cdots i_{k}$ to $i_{k+1}\cdots i_{m}$, it is possible to input an operator $\mathcal{O}$ with support on the input legs (as shown) and push the operator past the isometry, finding a new representation of the operator on the output legs.
Figure 1: Tensor-network diagrams depicting (a)-(b) the tensor $T_{i_{1}\cdots i_{m}}$ as it pertains to a quantum state, as well as an isometric map via the Choi-Jamiolkowski isomorphism [wilde2013quantum]. Here, red open legs represent logical input qubits, and black legs depict physical output qubits. (c)-(d) The action of tensor pushing an operator $\mathcal{O}$ through the input subsystems of the isometry in order to find an output-leg representation, which we name $\mathcal{O}^{\prime}$.
Figure 2(a) shows the support of a unitary logical operator $\mathcal{\bar{O}}$ on the output subsystems as $\mathcal{\bar{O}}=\mathcal{O}_{1}\otimes\cdots\otimes\mathcal{O}_{6}$ (in our specific case, for simplicity). In (b), we see that contracting two such tensors along their respective indices gives rise to a new conjoined tensor (in this case with two logical qubits). Finding the new logical operators for this new isometry is quite simple, and follows the tensor-matching formalism described in detail in [quantumlego, Farrelly_ltnc]. As shown in (c), the principle requirement is that support on contracted output indices must be identical up to complex conjugation; in the specific case of Figure 2(c), the requirements must be that $\mathcal{O}_{3}=\mathcal{O}^{\prime*}_{6}$ and $\mathcal{O}_{4}=\mathcal{O}^{\prime*}_{5}$.
Figure 2: (a) One may define an operator $\mathcal{\bar{O}}$ on an input leg to a tensor by its representation (in the present case $\mathcal{O}_{1}\otimes\cdots\otimes\mathcal{O}_{6}$) on the output legs. (b) Contracting indices together yields a new, conjoined tensor network. (c) Deducing the new logical operator is possible by keeping the original logical operators that satisfy the matching condition $\mathcal{O}_{3}=\mathcal{O}^{\prime*}_{6}$ and $\mathcal{O}_{4}=\mathcal{O}^{\prime*}_{5}$ [quantumlego, Farrelly_ltnc].
The concepts and operations explained above will be of paramount use in the software implementation, as we will see in Section III.
II-B Holographic Quantum Codes
Holographic quantum codes are a relatively new class of generalized concatenated codes [heterogeneous_holo_qrm, steinberg2024far, steinberg_2023]. These codes have been found to exhibit properties consistent with the AdS/CFT correspondence [harlow2015bulk, witten1998anti, maldacena1999large, gubser1998gauge], in addition to possessing interesting and potentially useful error-correction properties for practical quantum computing, such as: high, tunable rates [steinberg2024far, farrelly_parallel_decoding]; versatility in constructions involving several different seed codes [heterogeneous_holo_qrm]; distance scaling better than $\sqrt{n}$ [harris_int_opt]; and constructions with large fault-tolerant logical gate sets [heterogeneous_holo_qrm], among other useful properties [fan2024overcoming, junyu_thesis].
[…]
A holographic quantum code is built typically from either one or several seed codes, as shown in Figure 3. As hyperbolic tessellations allow for many different orientations and polygonal embeddings in the bulk subject to the constraint $\frac{1}{p}+\frac{1}{q}<\frac{1}{2}$, infinitely many variations of holographic quantum codes are possible to construct [heterogeneous_holo_qrm, ratcliffe1994foundations]. However, only homogeneous [harris_css_hqec, preskill_happy_code, farrelly_tn_codes, steinberg2024far] (that is, those holographic codes constructed from uniform hyperbolic tessellations with only one polygon featured, as shown in Figure 3(a)) and simple heterogeneous (with only alternating layers of two seed codes, as in Figure 3(b)) holographic codes [heterogeneous_holo_qrm] have so far been constructed. Growing a holographic code layerwise is done in two distinct methods, as shown in Figure 4: either using (a) edge inflation, in which every inflation step adds the polygons adjacent to the edges of the previous boundary, or (b) vertex inflation, in which each inflation step adds the polygons adjacent to the boundary vertices, resulting in a connected ring of polygons to be added. In the tensor-network scheme, subsequent layers must contract indices of the previous layer in order to adhere to the stringent constraints of a hyperbolic tessellation. More details on this topic and the motivation behind holographic codes can be found in [jahn_topical].
Figure 4: Methods by which one can perform holographic concatenation. Both (a) edge inflation and (b) vertex inflation of a tensor network on the $\{5,4\}$ face-centered hyperbolic tessellation [conformal_quasicrystals_holography, jahn2022tensor_qcft, jahn_topical]. In both cases, we show the first three inflation steps around a central pentagon.
III Software Design
LEGO_HQEC was developed in Python. For ease of use, the code is modularized into two parts: (a) the stabilizer and logical operator generator, and (b) the decoders. Both will be addressed in what follows.
III-A Stabilizer & Logical Operator Generation
Stabilizer and logical operator generation is carried out following several steps, which are depicted in Figure 5. Firstly, the object TensorNetwork is defined, with dependencies Tensor and TensorLeg as important classes; these are the components used to construct the holographic tensor network. Next, the Push function is defined; in this portion, each tensor has its own OperatorPush method on its local legs. Push then activates tensors one-by-one from the center to the boundary, and sequentially generates boundary representations of all locally-defined stabilizers and logical operators. Additionally, legs that have undergone an operator push will be blocked, in order to prevent operators from being accidentally pushed back in towards the center; in this way a locally-defined operator can be pushed to the boundary of the network. Finally, the ReadOut layer retrieves boundary stabilizers and logical operators using a backtracking method, and compiles them into the final boundary result; this backtracking method starts from the center of the tensor network, then uses a depth-first traversal to reach outer tensors, while recording operators on the legs of the outermost tensors at the boundary. With this method, it is ensured that boundary operators align with the geometric intuition expected from holographic codes. As a last step, a correctness check is performed by again traversing all internal legs and checking that all bulk-defined operators have been pushed to the boundary. More details on this portion of the software are available in [junyu_thesis].
Figure 5: Pictographic representation of the stabilizer / logical operator generator in LEGO_HQEC. TensorNetwork is a list containing all of the Tensor objects, whose attributes include the TensorLeg object. This list of tensors is then filtered into the Push function, which elementwise processes and pushes each operator from each tensor to the boundary, obtaining a boundary representation. Finally, a correctness checker (CorrectnessChecker) backtracks along the tensor network, checking to see that each contracted index from each tensor contains only the identity operator. The final boundary representation is read out with ReadOut.
III-B Decoders
Three decoders are included in LEGO_HQEC: An erasure decoder (Section III-B1) for the quantum erasure channel [wilde2013quantum, nielsen2010quantum] as well as two decoders for Pauli noise based on integer optimization and a tensor-network decoder (Section III-B2[farrelly_parallel_decoding, farrelly_tn_codes].
III-B1 Quantum Erasure Channel
[…]
$|,\mathcal{E}_{e}(\rho)=(1-\epsilon)\rho+\epsilon\text{Tr}\big[\rho\big]\ket{e}\bra{e}$~, (10)
[…]
LEGO_HQEC includes an erasure decoder based on Gaussian elimination (GE), as was utilized in [steinberg2024far, harris_css_hqec]. The general idea of the GE decoder involves stacking all of the stabilizers and logical operators into a large matrix, and then using the support of the erasure vector, for which entries are randomly generated per qubit with probability $p$, to filter out columns that have undergone an erasure. The algorithm then proceeds by executing Gaussian elimination to find the reduced-row-echelon form of the matrix; if it is found that both logical operators still have physical support on the remaining qubits, then the trial is counted as a success. The runtime complexity of the GE algorithm is known to scale as $\mathcal{O}(n(n-k)^{2})$. More details on the GE decoder can be found at [harris_css_hqec, steinberg2024far].
III-B2 Pauli Noise Channels
The prototypical Pauli noise channel is represented by the depolarizing noise channel $\mathcal{E}_{d}(\cdot)$ [nielsen2010quantum, wilde2013quantum], and takes the form
$\mathcal{E}_{d}(\rho)=(1-\delta)\rho+\frac{\delta}{3}\big(X\rho X+Y\rho Y+Z\rho Z\big)~,$ (11)
where $\delta$ is the probability of an error occurring, and $\{X,Y,Z\}$ represent the three Pauli operators. For Pauli noise, LEGO_HQEC employs two different types of decoders: one based on integer optimization [steinberg2024far, harris_int_opt], and another based on tensor-network decoding [fan2024overcoming, farrelly_tn_codes].
Integer optimization decoding works via a very similar technique to the GE decoder discussed in Section III-B1, but with some small changes: Firstly, as error locations are not known, one must first calculate the Moore-Penrose pseudoinverse [penrose1955generalized, moore1920reciprocal], which corresponds to the destabilizers (or pure errors) of the code [harris_int_opt, gottesman_improved_sim]. One then obtains a guessed error by multiplying the pseudoinverse matrix directly by the vector of the obtained syndrome. Finally, an integer optimization is utilized in order to deduce the lowest-weight pure-error correction [junyu_thesis, harris_int_opt]. It is known that integer-optimization decoding exhibits a runtime complexity of $\mathcal{O}(2^{n})$, which is dependent on the size of the quantum code. More details on integer-optimization decoding can be found in [junyu_thesis, steinberg2024far, harris_int_opt].
Tensor-network decoding, however, works by exploiting the tensor-network geometry and clever contraction sequences [farrelly_tn_codes, farrelly_parallel_decoding]. By setting the bond dimension of each tensor in the network to be four, Pauli errors $\{I,X,Y,Z\}$ can be easily accounted for. In this way, an error vector is initialized at random with some error probability $d$. These error probabilities are then utilized in the tensor network, and, upon contraction, yield a close approximation in the thermodynamic limit to maximum-likelihood decoding [chubb2021statistical]. It is known that tensor-network decoding exhibits a runtime complexity of $\mathcal{O}(n^{2.37})$ [le_gall_matrix_mult, bravyi_suchara] [^1] Although the runtime complexity above has been reported in several studies, we note that the overall runtime complexity for tensor-network decoding depends strongly on the tensor network’s geometry, as well as the contraction scheme (i.e., whether or not approximate contractions are used) and contraction ordering., and that multiple logical qubits can be decoded in parallel as $\mathcal{O}(nk)$, as long as the error rate stays below the critical error rate of the central logical qubit (i.e., the threshold of the central logical qubit) [farrelly_parallel_decoding]. More details on tensor-network decoding can be found in [joe_chubb, farrelly_tn_codes, farrelly_parallel_decoding].
IV Examples Using LEGO_HQEC
This section contains a step-by-step set of examples for utilizing LEGO_HQEC as well as a runtime analysis. We begin by initializing and testing a maximum-rate Harlow-Preskill-Pastawski-Yoshida (HaPPY) [preskill_happy_code] code under erasure noise, and then following up with an asymptotically zero-rate Steane code [harris_css_hqec] under Pauli depolarizing noise. The results of both decoder simulations are shown in Figure 6.
Figure 6: (a) Erasure decoder results for the maximum-rate HaPPY code (which obtains no threshold, as per the results in [preskill_happy_code]), which utilized the GE decoder (Section III-B1). (b) Pauli depolarizing noise results from the tensor-network decoder for an asymptotically zero-rate Steane code; the threshold of this code is found to be approximately $18.25\%$, in agreement with results from [harris_int_opt]. $1000$ Monte Carlo trials were utilized for the erasure and depolarizing noise simulations. Results are gathered per layer of concatenation $L$. The code snippets related to both (a) and (b) can be found in LABEL:code1 and LABEL:code2, respectively. Regarding computation time, for (a) the recovery rate curves for $L=1$ to $L=5$ took $14$ s, $15$ s, $18$ s, $536$ s, and $2204$ s respectively. In (b) the cases $L=0$ to $L=2$ took $15$ s, $194$ s, and $1497$ s, respectively (on an i9-14900K processor using 24 threads).
IV-A Usage
IV-A1 Maximum-Rate HaPPY Code under Erasure Noise
from LEGO_HQEC.QuDec.ErasureDecoder import calculate_recovery_rates_for_p_range
from LEGO_HQEC.OperatorPush.Presets.HaPPY_code import setup_max_rate_happy
from LEGO_HQEC.OperatorPush.PushingToolbox import batch_push
from LEGO_HQEC.QuDec.InputProcessor import extract_stabilizers_from_result_dict, extract_logicals_from_result_dict
from LEGO_HQEC.QuDec.OutputProcessor import save_results_to_csv
if __name__ == ’__main__’:
# Examine the HaPPY code with radii R=0, 1, 2, and 3, respectively.
for R in [1, 2, 3, 4, 5]:
# Obtain TN of the zero-rate HaPPY code with corresponding radius via preset.
tensor_list = setup_max_rate_happy(R=R)
results_dict = batch_push(tensor_list)
stabilizers = extract_stabilizers_from_result_dict(results_dict)
logical_zs, logical_xs = extract_logicals_from_result_dict(results_dict)
logical_operators = [logical_zs[0]] + [logical_xs[0]]
# Starting from an erasure error rate p of 0.0 to 1.0 with a step size of 0.05,
results = calculate_recovery_rates_for_p_range(n=1000, p_start=0.0, p_end=1.0, p_step=0.05, stabilizers=stabilizers, logical_operators=logical_operators)
save_results_to_csv(results, file_path=F’R{R}_rec.csv’)
Listing 1: Code snippet for initializing the stabilizers and logical operators of the maximum-rate HaPPY code, as well as for testing the first through fifth layers of concatenation under the erasure noise channel. Decoding was carried out using the GE decoder.
LABEL:code1 displays a code snippet for the construction and execution of the threshold calculation for the maximum-rate HaPPY code under erasure noise, utilizing the GE decoder discussed in Section III-B. In the example above, the commensurate libraries and dependencies are first imported (lines 1-5), and then the stabilizers and logical operators for each layer are generated (lines 11-13). Those objects are then fed into the erasure decoder per layer $L$, and the Monte Carlo trials commence layerwise (line 17). Finally, the results are written into a CSV file and saved (line 19).
IV-A2 Asymptotically Zero-Rate Steane Code under Depolarizing Noise
from LEGO_HQEC.QuDec.TN_decoder import tn_quantum_error_correction_decoder_multiprocess
from LEGO_HQEC.OperatorPush.Presets.Holographic_Steane_code import setup_heptagon_zero_rate_steane
from LEGO_HQEC.QuDec.OutputProcessor import save_results_to_csv
import numpy as np
if __name__ == ’__main__’:
rx = 1/3
rz = 1/3
ry = 1/3
for R in [0, 1, 2]:
tensor_list = setup_heptagon_zero_rate_steane(R=R)
p_depo_step = 0.05
p_depo_start = 0.0
p_depo_end = 0.5
succ_rates = []
for p_depo in np.arange(p_depo_start, p_depo_end + p_depo_step, p_depo_step):
print(f"Running at R = {R}, p_depo = {p_depo}")
succ_rate = tn_quantum_error_correction_decoder_multiprocess(tensor_list=tensor_list, p=p_depo, rx=rx, ry=ry, rz=rz, N=1000, n_process=8)
succ_rates.append((p_depo, succ_rate))
[…]
Listing 2: Code snippet for initializing the stabilizers and logical operators of the asymptotically zero-rate Steane code, as well as for testing the first three layers of concatenation under the Pauli depolarizing noise channel using the tensor-network decoder.
LABEL:code2 shows the code needed for initializing the tensor-network decoder for a depolarizing noise channel. In contrast to the previous case, the tensor network decoder requires a specification of the relative proportions of $X$, $Y$, and $Z$ errors (lines 7-9). However, the actual generation of the stabilizers and logical operators proceeds identically as for the HaPPY code (lines 14-17).
IV-B Runtime Performance
Here we report an example of the runtime performance expected of an asymptotically zero-rate HaPPY code for: the stabilizer and logical operator generation OperatorPush; the GE erasure decoder; the integer-optimization decoder; and the tensor-network decoder. All results are given as a function of layer $L$ and times are reported in seconds (s). For the decoders, the plotted results are per Monte Carlo trial, while the OperatorPush results denote the per-layer start-to-finish time for pushing all locally-defined stabilizers and logical operators to their respective boundary representations, and extracting them into a CSV file.
Figure 7: Time-scaling results per event for the OperatorPush function, as well as the three different decoders utilized in this software package. Due to limitations in computational power, we show the results regarding the exponentially-scaling integer-optimization algorithm for only the first two layers of the asymptotically zero-rate HaPPY code.
We were able to extract results for only the first two layers using the integer-optimization decoder, as we were limited by computational power. For all other decoders and the OperatorPush method, however, we were able to obtain time-scaling estimates for up to $L=4$, i.e., on the order of roughly several thousand boundary (physical) qubits, depending on the Schläfli symbols $\{p,q\}$ associated with the commensurate hyperbolic tiling.
V Results for Holographic Black Hole Codes
As a further demonstration of the range of our software tool’s usability, we investigate the threshold behavior of the black hole HaPPY code under the erasure channel, which was introduced in [heterogeneous_holo_qrm].
Figure 8: The black hole HaPPY code with five logical qubits (red legs), constructed by removing the central tensor from the zero-rate code of Figure 3(c). Here, we show the black hole code with (a) $L=2$ layers and (b) $L=3$ layers of edge inflation, corresponding to 95 and 375 physical qubits (outer boundary legs), respectively.
Instances of the code are constructed by removing the central tensor and then growing $L$ edge inflation layers of the hexagon code tiling (see Figure 8). Each instance encodes five logical qubits; stabilizers and logical operators are generated with OperatorPush, and decoding is performed using the Gaussian-elimination (GE) erasure decoder with $10^{5}$ Monte Carlo trials per $(p,L)$. Logical success probabilities are evaluated with respect to the simultaneous recoverability of all five logical qubits. We report a threshold $p_{\text{th}}=50\%$ from finite-size crossings (Figure 9). Consistent with our data, $p_{\text{th}}=50\%$ coincides with that of the zero-rate HaPPY code, while the black hole HaPPY code exhibits uniformly lower success probabilities; moreover, as $L$ increases the rate of the black hole HaPPY code approaches zero. Besides the black hole HaPPY code, we also constructed and tested the black hole holographic Steane code [harris2018calderbank] and the black hole holographic $[[6,1,3]]$ code [farrelly2021tensor], obtaining similar threshold results around $p_{\text{th}}=50\%$. These black hole codes are also asymptotically zero-rate; however, since they encode more logical qubits than their single-logical-qubit counterparts, their recovery curves lie slightly below while maintaining the same threshold.
Figure 9: Recovery probability of the black-hole HaPPY code under the erasure channel as a function of $P_{\text{erasure}}$ for $L\in\{1,\dots,4\}$. Each instance encodes five logical qubits, with success defined as simultaneous recovery of all. The threshold matches with that of the one-logical-qubit zero-rate HaPPY code, but with lower recovery probability. The black hole code is constructed by removing the tensor at layer $L=0$ from a regular holographic code (see Fig. 8); hence, the smallest code starts from $L=1$. For the simulations, $L\in\{1,2,3\}$ used 10,00010{,}000 Monte Carlo trials for each $P_{\text{erasure}}$ value, and $L=4$ used 1,0001{,}000 trials, with step size $\Delta p_{\text{erasure}}=0.01$. The computation times were $15$ s, $141$ s, $989$ s, and $3805$ s for $L\in\{1,\dots,4\}$, respectively (on an i9-14900K using 24 threads).
VI Conclusion
In this work, we have presented and demonstrated LEGO_HQEC, a new software tool for constructing and testing holographic quantum codes under typical noise channels of interest. Similar to other notable software packages for quantum error correction, fault tolerance, among others [qgym_feld, sarkar2024yaqq, wille_qldpc, gidney2021stim, qecsim, QiskitCommunity2017], our focus with LEGO_HQEC has been to facilitate open-source development and enable future scientific discovery in the realm of holographic quantum error correction. In the process, we have: developed and implemented an easy-to-use open-source software tool; demonstrated several examples of its capabilities; and have provided a basic framework upon which more detailed code studies and further discoveries in holographic quantum error correction can be leveraged.
LEGO_HQEC has already featured in several holographic code studies, including those exploring the novel Evenbly codes [steinberg2024far, steinberg_2023], analyses of asymptotically zero-rate holographic codes as compared to the hashing bound [fan2024overcoming, junyu_thesis], as well as for constructions of heterogeneous holographic codes [heterogeneous_holo_qrm], which are constructed from several seed codes and can circumvent the Eastin-Knill theorem regarding codes with universal logical gate sets [eastin2009restrictions, jochym2014using, chamberland2016thresholds].
As a proof of concept, we presented the black-hole HaPPY code and its threshold under the erasure noise channel. Our numerical analysis shows that, while its threshold under the erasure channel coincides with that of the one-logical-qubit zero-rate HaPPY code, the overall recovery performance is reduced for finite code instances: Though the asymptotic rate is zero for both codes, comparing instances at a fixed number of physical qubits leads to higher rate and lower logical recovery probability for the black-hole code.
There are many ideas for future modification of LEGO_HQEC. Firstly, an extension for considering any hyperbolic tiling parameterized by $\{p,q\}$ would be enormously helpful, as it has been shown that certain classes of holographic codes have vastly different threshold behavior as the tiling is altered [steinberg2024far]. Furthermore, recent work [heterogeneous_holo_qrm, junyu_thesis] shows that incorporating multiple seed codes can allow for potentially universal fault-tolerant logical operations [heterogeneous_holo_qrm]; tools to support such fault tolerance studies could be very useful. One can also consider implementing functionality for greater control over the encoded logical qubits in the bulk, as recent work shows that recovery rates of finite-rate holographic codes greatly depend not only on the number $k$ of encoded logical qubits, but critically on the positions of said logical qubits in the bulk [alex_logical_recovery]. Even more broadly, it is known that the qLEGO formalism applies to non-Abelian stabilizer codes, such as the XP stabilizer variety [xp_stab_quantum_cao, webster2022xp], or even non-stabilizer codes, such as those presented in [exotic_gatesets, lidar2013quantum, non_additive1, non_additive2, non_additive3]; one may consider incorporating functionality to accommodate such holographic code constructions in the near future. Moreover, the class of codes known as Evenbly codes, a derivate of the hyperinvariant tensor networks first introduced in [hyper_evenbly, steinberg2022conformal], have been extensively studied using LEGO_HQEC [steinberg2024far, steinberg_2023]; however, one major impediment was the inclusion of Hadamard gates on the contracted edges of the tensor network, thereby hindering decoding of Pauli channels with the tensor-network decoder. Allowing for such constraints, among others, remains within the future goals of this work. Finally, there are many potential additions purely from a software and usability standpoint, such as: adding functionality for generalized concatenated codes [knill2005quantum, gottesman1997stabilizer, lidar2013quantum]; visualization tools, such as those that have recently appeared in the software package PlanqTN [pato_planqTN]; or GPU acceleration tools; among other prospects.
VI-A Acknowledgements
We thank David Elkouss and Rob Harris for useful discussions. MS and SF acknowledge financial support from the Intel corporation. AJ is supported by the Einstein Research Unit “Perspectives of a quantum digital transformation” and Berlin Quantum. CC acknowledges support from the Commonwealth Cyber Initiative at Virginia Tech.
VI-B Author Contributions
JF and MS implemented the Gaussian elimination decoder. JF developed the integer-optimization decoder, tensor-network decoder, and the OperatorPush generator during his master thesis; the master thesis itself was supervised by MS and SF. AS provided assistance with software design. MS wrote the manuscript, with the hyperbolic tiling figures provided by AJ. CC, AJ, AS, and SF provided project guidance and supervision.